OpenStudy (anonymous):

write a specific formula to describe the variation: x varies directly with the cube root of y; x = 3 when y = 125

OpenStudy (anonymous):

x = k(y^1/3) 3 = k(125^1/3) 3 = 5k k = 3/5 Therefore, x = (3/5)(y^(1/3))

OpenStudy (anonymous):

Thats what I thought. This question has choices with it. x = - 5/ cbrt(y) x = 5 / cbrt(y) x = 15/cbrt(y) x = - 15 / cbrt(y) I get that the numerator is cbrt(y). But I got 3/5 for the numerator

OpenStudy (anonymous):

I do NOT see any error in what I did. If you see an error, please tell me where that is. I just dont see it.

OpenStudy (anonymous):

The cube root of y is represented by y^(1/3) varying directly with the cube root of y is k(y^(1/3))

OpenStudy (anonymous):

And I subsituted and solved correctly, from what I see.